# Efficient Adaptable High-throughput Streaming Aggregation Engines

### v3

@@ IV-B Experimental results – Group-by aggregation
This part of the evaluation explores the base pipeline as a standalone add-on to the existing unmodified small-scale sorter [15]. The supported operations per groupgroup ID in this implementation are: minimum, maximum, sum, count and distinct count. The engine accepts a data stream consisting of $(group,key)$ pairs. Each input tuple is 64-bit wide, 32-bit for each of the two fields. The pipeline throughput $P$ is 4 tuples per cycle.
@@
The C++ baseline is equivalent to the functionality of the design, i.e. first sorting the data and then performing the aggregation serially. The sorting function is std::sort, which is a high performing sorting baseline for general use, although not the absolute fastest [17]. The achieved speedup for $2^{16}$ tuples over the embedded processor varies between 22 to 28 according to the input distribution. This variation relates to the number of rows in the output. For instance, if all tuples have the same group ID, then the latency of interacting with the FPGA would be comparatively higher. This is because only a single tuple would be produced as an output.
@@
This speedup mostly comes from the sorter, as the studied CPU is already efficient for linearly skimming through a sorted stream while aggregating, also given its advantageous memory interfacing and speed. The speedup behaviour, such as for arbitrarily long streams, is inline with the sorter [15], hence not elaborating with comparisons for standalone group-by.
IV-C Experimental results – SWAG engine, without groups
@@
The SWAG engine is initially evaluated without group IDs, and a single stream of 32-bit integers is processed with high throughput. In this case, in order to saturate the 256-bit datapath, the generated aggregation pipeline and sorter-cardinality counter are for $P=8$, and without groupgroup fields. As with the operator type, the WS and WA values are also configured at runtime.
@@
This engine is compared against IBM’s reference implementations of multiple aggregation algorithms on the platform’s CPU [19] after porting it to ARM. These include the nbfinger, amta, two_stacks_lite, chunked_two_stacks_lite, daba_lite variations using the run_bulk_evict_insert.py experiment setup. The tested operations are the commonly supported sum and min. or max. for data in-order, though only sum is shown for brevity, as the performance remains relatively similar for these operators. Each kept datapoint is the minimum time from all algorithms, so that we compare with the best available software performance per window size (WS) and window advance (WA) combination.
@@
IV-D Experimental results – SWAG engine, with groups
@@ IV-E Higher-end baselines
Our system-level evaluation focuses on single-thread baselines to avoid introducing more variables. This is inline with reference implementations aiming at algorithmic comparison [19]. Specifically, SWAG exhibits a performance tradeoff between incremental computation and parallelism [20], making the selection of an oracle parallel implementation challenging. For the sake of completeness, a brief comparison is provided with an emulated parallel version on the same platform, and a high-end processor in figure 12. The parallel implementation is emulated by the single-thread version making fork() calls. Then, the median wall-time is divided by the number of “threads”. In this way, multi-processing inefficiencies such as memory pressure are partly reproduced, providing a lower bound for the obtained speedup.
@@
On the same platform, the FPGA speedup ranges from 13x to 148x over the parallel oracle version according to the input and operator. Over 7950X3D, the actual speedup for single-thread is up to 26x, while sum with a single group is slower due to its incremental nature and data localities. When using 32 threads, the performance is similar at up to 1.2x, while noting the same input/operator combination outlier. Note this processor has a more advanced data source (DDR5 vs LPDDR4) and process (5 vs 16 nm). The FPGA power is also notably lower at 3.34W over 45W for this workload (latter influenced by DVFS), highlighting its appropriateness for edge applications, which are prominent with SWAG [1].
@@ IV-F Resource utilisation
TABLE I: Vivado resource utilisation report summary
@@
With respect to the group-by aggregate engine, we can see that there is about a 2.5K overhead on the number of Lookup-tables (LUTs) when becoming an add-on to an existing unmodified sorter. All engine designs here use a register per pipeline stage, which also group-by introduced a 2.4K overhead on flip-flop (FF) registers for this engine.
@@ IV-G Scalability
The proposed engines are scalable to high degrees of parallelism $P$ and are able to saturate the bandwidth for current and future systems with a wide datapath. The total amount of data that can be produced or consumed by the engine per cycle is $P$ times the data width. An exploration of the impact of $P$ and the data width on the resources is conducted through out-of-context synthesis. The tuple (data) width is halved and used by the group ID and key fields for simplicity. The synthesis is done with Yosys [21] for a more generic but modern LUT6 architecture, which can serve as high-level guidelines for the resource utilisation expectations for a variety of platforms.
@@
Figure 13 summarises the results of this exploration for the SWAG engine without the accompanying sliding window buffering and sorter. The sliding window buffering logic is simplistic, hence it is omitted from the scalability study, as it mainly consists of block RAM. Since our modifications to the sorter are relatively lightweight, the readers can also refer to the sorter’s original description on its scalability [15]. For the ranges of interest, the registers and LUTs scale almost linearly with $P$ and the tuple width. The fastest growing component is the reverse butterfly network, with a complexity of $\Theta(P\log_{2}(P))$ switches (similar for the prefix scans), though the slight upward trend is not observed in these ranges yet. Our use case has yielded a low enough register utilisation for our purposes, but future work can also focus on retiming to considerably reduce the register utilisation when necessary.
@@
This scalability study focuses on the aggregation part for SWAG with groups without loss of generality, since it is the most feature-rich version of the base pipeline. The obtained curves for the other pipelines are very similar, but they are omitted for brevity.
@@ IV-H Throughput and latency
Aggregation engines are often found as streaming accelerators, hence it can be a requirement to be able to cope with the line rate of the system, especially when the stream is continuous. As a non-blocking pipeline, the main aggregation pipeline supports the full throughput of $P$ elements per cycle without any backpressure for tasks such as group-by. Its pipeline latency of $1+2\log_{2}P$ cycles is negligible for streaming data, which could be minimised further with register retiming.
When it is extended to implement the SWAG engine, it becomes a blocking design (i.e. its “ready” signal or equivalent is not always high). The amount of backpressure it applies on its producer mostly relates to the window advance (WA) in relation to its window size (WS). This is because, when $\mathrm{\textit{WA}}<\mathrm{\textit{WS}}$ it has the potential to produce more output than input [22]. Specifically, assuming that other than the sliding window buffering there is no other backpressure from the rest of the engine or system, the highest throughput that can be achieved by an ideal implementation is $P\mathrm{\times\textit{WS}}/\mathrm{\textit{WA}}$, where $P$ is the degree of parallelism.
@@
Figure 14 provides simulation results using Icarus Verilog [23] that show the input rate by counting the number of input tuples per cycle for a relatively long input size and $P=4$. On the left plot, it can be observed that the input rate peaks near 4, especially when $\mathrm{\textit{WS}}=\mathrm{\textit{WA}}=k=128$. This is since the sorter processes the data as they arrive, there is no buffer “replay” for $\mathrm{\textit{WS}}=\mathrm{\textit{WA}}$, and any other buffering within the design such as the PMT is comparatively diminishing. After $\mathrm{\textit{WS}}=256$, the maximum input rate is up to half of the available, since the sorter needs an internal second pass for the merging phase. On the right plot, the numbers are divided by $\mathrm{P\times\textit{WS}}/\mathrm{\textit{WA}}$ to uncover any inefficiencies, with the main observation being the halving of the throughput when $\mathrm{\textit{WS}}>k$.
The plots of figure 14 have a triangular shape, as the supported window advance cannot be greater than the window size, i.e. $\mathrm{\textit{WA}}\leq\mathrm{\textit{WS}}$. At $\mathrm{\textit{WS}}=256$, there are no results due to a limitation within the sorter (merging at least $P$ lists of $k$ elements). If the missing $\textit{WS}=256$ mode is needed, a build with $k=64$ or $256$ would suffice. Alternatively, this inherited limitation can be addressed at its source [15], to enable splitting the sorted sublists into smaller pieces.
@@ V Related work
There is a moderate amount of research on database acceleration focusing on analytics [24]. Similar to earlier works that accelerated joins [25, 2], there are tradeoffs when selecting between a sort and an indexing-based solution [26, 27], each with different adaptation challenges [28, 29]. For instance, even with a sorter-based pipeline, supporting arbitrary data in a join accelerator caused stalls and increased its memory requirements [25]. This relates to the existence of duplicates of unknown cardinality, which is somewhat reminiscent of median support, which affects the complexity of SWAG designs.
@@
DFE [30] StreamZip [7] Presented engine implementations Mueller et al. [31] Oge et al. [32]
@@ V-A Group-by aggregation
Group-by aggregate operators have usually been a part of larger analytics designs [33]. This is due to the bandwidth limitations [29], as with co-processors’ reduced performance when it comes to accessing main memory. This makes stand-alone solutions more challenging to become practical.
A group-by-aggregate-focused work demonstrated a speedup of 10x over the CPU [27], but used multi-tasking, hashes and off-chip memory on a discontinued platform. A recent group-by-aggregator uses sorting to accelerate these operations [34], though it implements an SQL-based data management system as a whole, which is still inline with the observation on the scarcity of out-of-context group-by-aggregate accelerators.
Despite this, the acceleration of specific group-by queries is useful in certain industrial contexts, such as distinct count per group for customer analytics [15]. This operator is supported by the presented engine, but can also be implemented by combining two PRRAs for intermediate sparsity removal, as shown in figure 16.
Figure 16: Modular design of the distinct count query (right) [15] with example data in orange, based on the PRRA implementation (left) [4].
When we compare the engine’s utilisation (2.55K LUT, 2.42K FF, see table I) to the unmodified PRRA (1.26K LUT, 2.29K FF [4]), it is easy to conclude that it uses at least 1.9x fewer FFs than any solution with two PRRAs, and similar LUTs. This can be explained with the help of a simple model: when we abstract the node complexity to 1 for the butterfly network switches and all other entities, the complexity of the example PRRA-based engine of figure 16 decreases from $3P+2\times\mathrm{PRRA}$ to $2P+1\times\mathrm{PRRA}=2P\times\log_{2}(P)+P+1$ entities. This is while supporting multiple operators reconfigurable at runtime. In this instance, our proposed group-by engine reduces the best known resource complexity for performing distinct count per group, while being highly adaptable for other such advanced operators.
@@ V-B SWAG engines
Starting from the left of table II, the comparison begins with SWAG with groups. State-of-the-art FPGA solutions provide a notable speedup over prior work [30], but still operate on up to a single tuple per cycle [7, 35]. They often involve the implementation of a multi-level memory hierarchy based on off-chip memories, which is reflected in the complexity of the design. In contrast, our proposed engine achieves to be evaluated on a resource-constrained embedded platform (no loss of generality to high-end FPGAs).
An important difference between the presented SWAG engine and the selection of competing engines for SWAG with groups is that there is a single sliding window, as opposed to a window per group (elaborated in section VI). In this case, our design uses no DRAM and a fraction of the BRAM at 3.3 versus 22.4 Mbits for the most lightweight of the alternatives [30]. The competing engines have a target frequency of 150 MHz, but at 125 MHz internally, it is able to saturate a line rate over 3 times higher. In terms of throughput, the competing variations in the table are essentially exchanging resources and functionality for performance (not shown). For example, DFE utilises around $1/7$ of the input rate, even when WS==WA. The compression in StreamZip [7] causes the throughput to scale differently according to the relationship between WS and WA, which can be beneficial in certain applications with a low WA.
@@
When it comes to SWAG without groups, Mueller et al. propose a sliding median operator, but was more of a proof of a generalisable concept, also given the state of the technology of the day [31]. The sliding window size is fixed to only 8, hence the reduced need for BRAM and DRAM. The median calculation is based on shift registers and a sorting network, which, even with newer technology, do not have the scalability of more sophisticated designs. For instance, the merge-capable linear sorters [15] enable a practical use with WS of thousands of elements in the presented engines. It operates at a single input per cycle, though parallelism is demonstrated for multiple workers operating on independent streams. A single worker has a throughput of 100 MT/s, while our approach reaches 1 GT/s.
@@
Overall, our SWAG engine’s contributions stand between SWAG with groups and without groups including, enabling big data processing with more advanced operators for simple devices. It is worth noting that the comparison is, hence, relatively diverse, as it features partly different systems and operator combinations, and further customisation is possible.

### v4

@@ IV-B Experimental results – Enthuse, group-by aggregation
This part of the evaluation explores Enthuse as a standalone add-on to the existing unmodified small-scale sorter [fsorter]. The supported operations per groupgroup ID in this implementation are: minimum, maximum, sum, count and distinct count. The engine accepts a data stream consisting of $(group,key)$ pairs. Each input tuple is 64-bit wide, 32-bit for each of the two fields. The pipeline throughput $P$ is 4 tuples per cycle.
@@
The C++ baseline is equivalent to the functionality of the design, i.e. first sorting the data and then performing the aggregation serially. The sorting function is std::sort, which is a high performing sorting baseline for general use, although not the absolute fastest [flimsj]. The achieved speedup for $2^{16}$ tuples over the embedded processor varies between 22 to 28 according to the input distribution. This variation relates to the number of rows in the output. For instance, if all tuples have the same group ID, then the latency of interacting with the FPGA would be comparatively higher. This is because only a single tuple would be produced as an output.
@@
This speedup mostly comes from the sorter, as the studied CPU is already efficient for linearly skimming through a sorted stream while aggregating, also given its advantageous memory interfacing and speed. The speedup behaviour, such as for arbitrarily long streams, is inline with the sorter [fsorter], hence not elaborating with comparisons for standalone group-by.
IV-C Experimental results – EnthuseSWAG, without groups
@@
The engine is initially evaluated for SWAG without group IDs, and a single stream of 32-bit integers is processed with high throughput. In this case, in order to saturate the 256-bit datapath, the generated aggregation pipeline and sorter-cardinality counter are for $P=8$, and without groupgroup fields. As with the operator type, the WS and WA values are also configured at runtime.
@@
EnthuseSWAG is compared against IBM’s reference implementations of multiple aggregation algorithms on the platform’s CPU [ibm] after porting it to ARM. These include the nbfinger, amta, two_stacks_lite, chunked_two_stacks_lite, daba_lite variations using the run_bulk_evict_insert.py experiment setup. The tested operations are the commonly supported sum and min. or max. for data in-order, though only sum is shown for brevity, as the performance remains relatively similar for these operators. Each kept datapoint is the minimum time from all algorithms, so that we compare with the best available software performance per window size (WS) and window advance (WA) combination.
@@
IV-D Experimental results – EnthuseSWAG, with groups
@@ IV-E Higher-end baselines
Our system-level evaluation focuses on single-thread baselines to avoid introducing more variables. This is inline with reference implementations aiming at algorithmic comparison [ibm]. Specifically, SWAG exhibits a performance tradeoff between incremental computation and parallelism [theodorakis2020lightsaber], making the selection of an oracle parallel implementation challenging. For the sake of completeness, a brief comparison is provided with an emulated parallel version on the same platform, and a high-end processor in figure 12. The parallel implementation is emulated by the single-thread version making fork() calls. Then, the median wall-time is divided by the number of “threads”. In this way, multi-processing inefficiencies such as memory pressure are partly reproduced, providing a lower bound for the obtained speedup.
@@
On the same platform, the FPGA speedup ranges from 13x to 148x over the parallel oracle version according to the input and operator. Over 7950X3D, the actual speedup for single-thread is up to 26x, while sum with a single group is slower due to its incremental nature and data localities. When using 32 threads, the performance is similar at up to 1.2x, while noting the same input/operator combination outlier. Note this processor has a more advanced data source (DDR5 vs LPDDR4) and process (5 vs 16 nm). The FPGA power is also notably lower at 3.34W over 45W for this workload (latter influenced by DVFS), highlighting its appropriateness for edge applications, which are prominent with SWAG [Tangwongsan2018].
@@ IV-F Resource utilisation
TABLE I: Vivado resource utilisation summary of Enthuse
@@
With respect to the group-by aggregate engine (Enthuse), we can see that there is about a 2.5K overhead on the number of Lookup-tables (LUTs) when becoming an add-on to an existing unmodified sorter. All engine designs here use a register per pipeline stage, which also group-by introduced a 2.4K overhead on flip-flop (FF) registers for this engine.
@@ IV-G Scalability
The proposed engines are scalable to high degrees of parallelism $P$ and are able to saturate the bandwidth for current and future systems with a wide datapath. The total amount of data that can be produced or consumed by the engine per cycle is $P$ times the data width. An exploration of the impact of $P$ and the data width on the resources is conducted through out-of-context synthesis. The tuple (data) width is halved and used by the group ID and key fields for simplicity. The synthesis is done with Yosys [wolf2016yosys] for a more generic but modern LUT6 architecture, which can serve as high-level guidelines for the resource utilisation expectations for a variety of platforms.
@@
Figure 13 summarises the results of this exploration for EnthuseSWAG without the accompanying sliding window buffering and sorter. The sliding window buffering logic is simplistic, hence it is omitted from the scalability study, as it mainly consists of block RAM. Since our modifications to the sorter are relatively lightweight, the readers can also refer to the sorter’s original description on its scalability [fsorter]. For the ranges of interest, the registers and LUTs scale almost linearly with $P$ and the tuple width. The fastest growing component is the reverse butterfly network, with a complexity of $\Theta(P\log_{2}(P))$ switches (similar for the prefix scans), though the slight upward trend is not observed in these ranges yet. Our use case has yielded a low enough register utilisation for our purposes, but future work can also focus on retiming to considerably reduce the register utilisation when necessary.
@@
This scalability study focuses on the aggregation part for EnthuseSWAG with groups without loss of generality, since it is the most feature-rich version of the Enthuse pipeline. The obtained curves for the other pipelines are very similar, but they are omitted for brevity.
@@ IV-H Throughput and latency
Aggregation engines are often found as streaming accelerators, hence it can be a requirement to be able to cope with the line rate of the system, especially when the stream is continuous. As a non-blocking pipeline, Enthuse supports the full throughput of $P$ elements per cycle without any backpressure for tasks such as group-by. Its pipeline latency of $1+2\log_{2}P$ cycles is negligible for streaming data, which could be minimised further with register retiming.
When Enthuse is extended to implement EnthuseSWAG, it becomes a blocking design (i.e. its “ready” signal or equivalent is not always high). The amount of backpressure it applies on its producer mostly relates to the window advance (WA) in relation to its window size (WS). This is because, when $\mathrm{\textit{WA}}<\mathrm{\textit{WS}}$ it has the potential to produce more output than input [10.1145/3465998.3466014]. Specifically, assuming that other than the sliding window buffering there is no other backpressure from the rest of the engine or system, the highest throughput that can be achieved by an ideal implementation is $P\mathrm{\times\textit{WS}}/\mathrm{\textit{WA}}$, where $P$ is the degree of parallelism.
@@
Figure 14 provides simulation results using Icarus Verilog [williams2002icarus] that show the input rate by counting the number of input tuples per cycle for a relatively long input size and $P=4$. On the left plot, it can be observed that the input rate peaks near 4, especially when $\mathrm{\textit{WS}}=\mathrm{\textit{WA}}=k=128$. This is since the sorter processes the data as they arrive, there is no buffer “replay” for $\mathrm{\textit{WS}}=\mathrm{\textit{WA}}$, and any other buffering within the design such as the PMT is comparatively diminishing. After $\mathrm{\textit{WS}}=256$, the maximum input rate is up to half of the available, since the sorter needs an internal second pass for the merging phase. On the right plot, the numbers are divided by $\mathrm{P\times\textit{WS}}/\mathrm{\textit{WA}}$ to uncover any inefficiencies, with the main observation being the halving of the throughput when $\mathrm{\textit{WS}}>k$.
The plots of figure 14 have a triangular shape, as the supported window advance cannot be greater than the window size, i.e. $\mathrm{\textit{WA}}\leq\mathrm{\textit{WS}}$. At $\mathrm{\textit{WS}}=256$, there are no results due to a limitation within the sorter (merging at least $P$ lists of $k$ elements). If the missing $\textit{WS}=256$ mode is needed, a build with $k=64$ or $256$ would suffice. Alternatively, this inherited limitation can be addressed at its source [fsorter], to enable splitting the sorted sublists into smaller pieces.
@@ V Related work
There is a moderate amount of research on database acceleration focusing on analytics [fang2020memory]. Similar to earlier works that accelerated joins [casper2014hardware, owaida2017centaur], there are tradeoffs when selecting between a sort and an indexing-based solution [moghaddamfar2023study, absalyamov2016fpga], each with different adaptation challenges [matas2022automated, eryilmaz2021fpga]. For instance, even with a sorter-based pipeline, supporting arbitrary data in a join accelerator caused stalls and increased its memory requirements [casper2014hardware]. This relates to the existence of duplicates of unknown cardinality, which is somewhat reminiscent of median support, which affects the complexity of SWAG designs.
@@
DFE [8977927] StreamZip [geethakumari2023stream] EnthuseSWAG implementations Mueller et al. [mueller2009data] Oge et al. [6726885]
@@ V-A Group-by aggregation
Group-by aggregate operators have usually been a part of larger analytics designs [lisa2018column]. This is due to the bandwidth limitations [eryilmaz2021fpga], as with co-processors’ reduced performance when it comes to accessing main memory. This makes stand-alone solutions more challenging to become practical.
A group-by aggregate-focused work demonstrated a speedup of 10x over the CPU [absalyamov2016fpga], but used multi-tasking, hashes and off-chip memory on a discontinued platform. A recent group-by-aggregator uses sorting to accelerate these operations [xue2024fpga], though it implements an SQL-based data management system as a whole, which is still inline with the observation on the scarcity of out-of-context group-by-aggregate accelerators.
Despite this, the acceleration of specific group-by queries is useful in certain industrial contexts, such as distinct count per group for customer analytics [fsorter]. This operator is supported by the presented engine, but can also be implemented by combining two PRRAs for intermediate sparsity removal, as shown in figure 16.
Figure 16: Modular design of the distinct count query (right) [fsorter] with example data in orange, based on the PRRA implementation (left) [prra].
When we compare the engine’s utilisation (2.55K LUT, 2.42K FF, see table I) to the unmodified PRRA (1.26K LUT, 2.29K FF [prra]), it is easy to conclude that it uses at least 1.9x fewer FFs than any solution with two PRRAs, and similar LUTs. This can be explained with the help of a simple model: when we abstract the node complexity to 1 for the butterfly network switches and all other entities, the complexity of the example PRRA-based engine of figure 16 decreases from $3P+2\times\mathrm{PRRA}$ to $2P+1\times\mathrm{PRRA}=2P\times\log_{2}(P)+P+1$ entities. This is while supporting multiple operators reconfigurable at runtime. In this instance, our proposed group-by engine reduces the best known resource complexity for performing distinct count per group, while being highly adaptable for other such advanced operators.
@@ V-B SWAG engines
Starting from the left of table II, the comparison begins with SWAG with groups. State-of-the-art FPGA solutions provide a notable speedup over prior work [8977927], but still operate on up to a single tuple per cycle [geethakumari2023stream, gong2021shuntflowplus]. They often involve the implementation of a multi-level memory hierarchy based on off-chip memories, which is reflected in the complexity of the design. In contrast, our proposed engine achieves to be evaluated on a resource-constrained embedded platform (no loss of generality to high-end FPGAs).
An important difference between the presented SWAG engine and the selection of competing engines for SWAG with groups is that there is a single sliding window, as opposed to a window per group (elaborated in section VI). In this case, our design uses no DRAM and a fraction of the BRAM at 3.3 versus 22.4 Mbits for the most lightweight of the alternatives [8977927]. The competing engines have a target frequency of 150 MHz, but at 125 MHz internally, it is able to saturate a line rate over 3 times higher. In terms of throughput, the competing variations in the table are essentially exchanging resources and functionality for performance (not shown). For example, DFE utilises around $1/7$ of the input rate, even when WS==WA. The compression in StreamZip [geethakumari2023stream] causes the throughput to scale differently according to the relationship between WS and WA, which can be beneficial in certain applications with a low WA.
@@
When it comes to SWAG without groups, Mueller et al. propose a sliding median operator, but was more of a proof of a generalisable concept, also given the state of the technology of the day [mueller2009data]. The sliding window size is fixed to only 8, hence the reduced need for BRAM and DRAM. The median calculation is based on shift registers and a sorting network, which, even with newer technology, do not have the scalability of more sophisticated designs. For instance, the merge-capable linear sorters [fsorter] enable a practical use with WS of thousands of elements in the presented engines. It operates at a single input per cycle, though parallelism is demonstrated for multiple workers operating on independent streams. A single worker has a throughput of 100 MT/s, while our approach reaches 1 GT/s.
@@
Overall, EnthuseSWAG’s contributions stand between SWAG with groups and without groups including, enabling big data processing with more advanced operators for simple devices. It is worth noting that the comparison is, hence, relatively diverse, as it features partly different systems and operator combinations, and further customisation is possible.